# DIVER-1: Deep Integration of Vast Electrophysiological Recordings at Scale

### v1

@@ D.4 Analysis of Joint Modality (intracranial/scalp-EEG) Pretraining
Table 21: Ablation on pretraining data composition for EEG downstream tasks. We evaluated how different pretraining dataset combinations affect performance on EEG downstream tasks. We compared models pretrained on: TUEG (obeid2016tueg)-only (largest single EEG dataset), TUEG+iEEG, all EEG datasets, and EEG+iEEG (all available data). All models use DIVERSmall/E/1s architecture with $d_{\text{model}}=512$, pretrained for 2 epochs. The model trained exclusively on curated EEG datasets achieves the best performance, outperforming models that include iEEG data.
Pretraining Data FACED (9-class) PhysioNet-MI (4-class)
ACC kappa F1 ACC kappa F1
TUEG-only $0.519\pm 0.004$ $0.456\pm 0.005$ $0.518\pm 0.005$ $0.623\pm 0.010$ $0.497\pm 0.013$ $0.625\pm 0.009$
TUEG + iEEG $0.461\pm 0.011$ $0.394\pm 0.013$ $0.471\pm 0.012$ $0.599\pm 0.008$ $0.465\pm 0.010$ $0.600\pm 0.008$
@@
EEG $\mathbf{0.570\pm 0.009}$ $\mathbf{0.515\pm 0.010}$ $\mathbf{0.579\pm 0.008}$ $\mathbf{0.644\pm 0.005}$ $\mathbf{0.526\pm 0.006}$ $\mathbf{0.646\pm 0.005}$
@@
Head Depth FACED (9-class) PhysioNet-MI (4-class)
ACC kappa F1 ACC kappa F1
Linear $0.523\pm 0.018$ $0.461\pm 0.020$ $0.521\pm 0.018$ $0.653\pm 0.013$ $0.538\pm 0.017$ $0.654\pm 0.013$
2-Layer $0.584\pm 0.005$ $0.530\pm 0.005$ $0.584\pm 0.005$ $0.674\pm 0.003$ $0.565\pm 0.004$ $0.676\pm 0.003$
3-Layer $0.601\pm 0.008$ $0.550\pm 0.009$ $0.607\pm 0.009$ $\mathbf{0.676\pm 0.003}$ $\mathbf{0.567\pm 0.004}$ $\mathbf{0.678\pm 0.004}$
4-Layer $\mathbf{0.603\pm 0.007}$ $\mathbf{0.552\pm 0.008}$ $\mathbf{0.609\pm 0.007}$ $0.672\pm 0.008$ $0.573\pm 0.011$ $0.674\pm 0.008$
@@ D.5 Interpretation Results
Figure 13 shows a visualization of representation analysis on downstream datasets using UMAP (mcinnes2018umap). We examine the embeddings obtained from the pretrained $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ model on the test set of speech, onset, and volume tasks in neuroprobe dataset without any finetuning. The results indicate that DIVER learned meaningful iEEG representations from pretraining, thereby capturing label-relevant structure in downstream datasets even in the absence of finetuning.
@@ E.1 Overview:Impact of Finetuning Methodology on EEG Downstream Task Performance
We conduct comprehensive experiments to investigate the impact of different finetuning methods on downstream task performance. Our analysis includes: (1) systematic exploration of MLP classifier depth for DIVER(appendix subsection E.2), (2) reproduction of CBraMod (wang2024cbramod) performance using publicly available code and weights(appendix subsection E.3), and (3) comparative analysis of various finetuning configurations across both models E.4. These experiments reveal that finetuning methodology significantly affects model performance, and optimal configurations vary across tasks and models. Importantly, when comparing DIVER and CBraMod under identical finetuning configurations (one-to-one comparison), DIVER demonstrates competitive or superior performance to CBraMod across multiple tasks, achieving overall state-of-the-art results despite CBraMod’s higher reported in-paper performance on some tasks.
@@ E.2 Impact of MLP Classifier Depth on DIVER Performance
Initially, we employed a linear classifier head for finetuning but observed suboptimal performance. Upon examining the publicly available code of CBraMod (wang2024cbramod), the previous state-of-the-art model, we found that it uses an MLP classifier head. To ensure fair comparison, we replaced the linear head with an MLP classifier, which resulted in substantial performance improvements for DIVER-1. This motivated us to conduct systematic experiments investigating how MLP depth affects performance. Table 22 compares the performance across the depths of MLP classifier for downstream EEG tasks. The original 3-layer MLP classifier is varied between 1 to 5 layers. For both FACED and PhysioNet-MI tasks, performance improved with the increase of MLP depth, achieving peak performance at 4-layer for FACED (balanced accuracy of 0.603) and 3-layer for PhysioNet-MI (balanced accuracy of 0.676). Beyond the optimal depth, we observed performance saturation or slight degradation, particularly notable in the 5-layer MLP for both datasets. These results indicate that a moderate depth (3-4 layers) suffices model’s effectiveness across different EEG downstream tasks.
@@ E.3 Challenges in Reproducing CBraMod Baseline Performance
The experimental results revealed substantial performance gaps on several tasks (Table 23). On MentalArithmetic(zyma2019electroencephalograms), reproduced accuracy (0.619) fell short of reported performance (0.726); on TUEV(obeid2016tueg), accuracy decreased from 0.667 to 0.605; on Mumtaz2016(zyma2019electroencephalograms), from 0.956 to 0.882. Only FACED(chen2023faced) and PhysioNet-MI(goldberger2000physionet; schalk2004bci2000) showed relatively successful reproduction. Notably, similar reproduction difficulties with CBraMod have been reported in recent work (wang2025eeg), which also observed performance gaps between reported and reproduced results on certain tasks.
@@
Table 23: Comparison of DIVER-1 EEG model with CBraMod. CBraMod (in paper) refers to the performance reported in the original CBraMod paper, while CBraMod (reproduction) represents our reproduction using the default configuration from their publicly released code and weights.
@@ E.4 Task-Specific Optimal Finetuning Configurations for CBraMod and DIVER
Table 24: One-to-One Comparison between CBraMod and DIVER using fixed finetuning method. Both models evaluated with identical finetuning configurations. CBraMod uses publicly released weights and code. DIVER model uses 12 layers with $d_{\text{model}}=512$ and were pretrained for 32 epochs. Results are reported as mean $\pm$ standard deviation across 5 random seeds.
@@ F.1 Pretraining Dataset Description
  • AJILE12 (Anootated Joints in Long-term Electrocorticography) (peterson2022ajile12): An ECoG dataset from 12 epilepsy patients, recorded semi-continuously over 55 days. Signals were collected from $\geq$ 64 electrodes at 1 kHz sampling rate and paired with synchronized video-vased 3D human pose estimation and annotated wrist-movement events.
@@
  • TUEG (Temple University Hospital EEG Corpus) (obeid2016tueg): A large-scale clinical EEG dataset comprising 16,986 recording sessions from 10,874 subjects with heterogeneous diagnoses. EEG signals were recorded using 20–31+ channels, predominanty at sampling rate between 250–512 Hz, and are linked with de-identified clinical reports.
  • HBN-EEG (Healthy Brain Network) (shirazi2024hbn): An developmental EEG dataset from 2,782 participants aged 5–21. Each participant underwent approximately 60 minutes of high-density (128-channel) EEG and eye-tracking recordings across six distinct tasks, including resting-state and movie watching.
  • NCHSDB (Nationwide Children’s Hospital Sleep DataBank) (lee2022nsrr): An pediatric sleep EEG dataset of 3,673 patients. Each record includes 8–12 hours of EEG data (26–29 channels, sampled at 256–512 Hz) manually scored for sleep stages and events. While the dataset contains multimodal PSG signals (EOG, EMG, ECG, respiration, etc.), we used only the EEG channels.
  • PEERS (Penn Electrophysiology of Encoding and Retrieval Study) (peers): An EEG dataset from 364 subjects who participated in multiple sessions of free recall, recognitionm and distractor tasks. EEG signals were recorded with 125 channels at 500 Hz sampling rate.

### v2

@@ D.4 Analysis of Joint Modality (intracranial/scalp-EEG) Pretraining
To further analyze how the composition of pretraining data influences EEG downstream performance, we conducted ablation studies across different EEG and iEEG dataset combinations, as shown in Table 21. With only two pretraining epochs, models pretrained exclusively on curated EEG datasets consistently outperformed those incorporating iEEG data across both FACED and PhysioNet-MI benchmarks. Even within the EEG modality, combining multiple datasets outperformed using the largest single dataset alone (TUEG-only), highlighting the benefits of dataset diversity.
This trend, however, reverses with extended pretraining. As shown in Table 22, while EEG-only pretraining yields superior performance in epoch2 and epoch8, models jointly pretrained on EEG and iEEG eventually outperform EEG-only counterparts given sufficient training time(epoch16). This suggests that modality heterogeneity acts as an initial impediment that the model gradually overcomes, ultimately enabling it to leverage complementary information across modalities. Based on these findings, we selected the EEG+iEEG jointly pretrained model as our final configuration for all subsequent EEG downstream evaluations.
@@
Table 21: Ablation on pretraining data composition for EEG downstream tasks. We evaluated how different pretraining dataset combinations affect performance on EEG downstream tasks. We compared models pretrained on: TUEG (Obeid and Picone, 2016)-only (largest single EEG dataset), TUEG+iEEG, all EEG datasets, and EEG+iEEG (all available data). All models use DIVERSmall/E/1s architecture with $d_{\text{model}}=512$, pretrained for 2 epochs. The model trained exclusively on curated EEG datasets achieves the best performance, outperforming models that include iEEG data.
[…]
black Pretraining Data FACED (9-class) PhysioNet-MI (4-class) ACC kappa F1 ACC kappa F1 TUEG-only $0.519\pm 0.004$ $0.456\pm 0.005$ $0.518\pm 0.005$ $0.623\pm 0.010$ $0.497\pm 0.013$ $0.625\pm 0.009$ TUEG + iEEG $0.461\pm 0.011$ $0.394\pm 0.013$ $0.471\pm 0.012$ $0.599\pm 0.008$ $0.465\pm 0.010$ $0.600\pm 0.008$ EEG + iEEG $0.540\pm 0.013$ $0.482\pm 0.015$ $0.550\pm 0.014$ $0.638\pm 0.005$ $0.517\pm 0.006$ $0.639\pm 0.004$ EEG $\mathbf{0.570\pm 0.009}$ $\mathbf{0.515\pm 0.010}$ $\mathbf{0.579\pm 0.008}$ $\mathbf{0.644\pm 0.005}$ $\mathbf{0.526\pm 0.006}$ $\mathbf{0.646\pm 0.005}$
Table 22: Ablation on pretraining data composition across pretraining epochs. We compare EEG-only vs EEG+iEEG pretraining at different pretraining epochs. All models use DIVERSmall/E/1s architecture with $d_{\text{model}}=512$. Notably, with longer pretraining, EEG+iEEG consistently outperforms EEG-only across downstream tasks.
[…]
black
FACED (9-class) PhysioNet-MI (4-class)
Epochs Pretraining Data ACC kappa F1 ACC kappa F1
2 EEG-only $\mathbf{0.570\pm 0.009}$ $\mathbf{0.515\pm 0.010}$ $\mathbf{0.579\pm 0.008}$ $\mathbf{0.644\pm 0.005}$ $\mathbf{0.526\pm 0.006}$ $\mathbf{0.646\pm 0.005}$
@@
8 EEG-only $\mathbf{0.594\pm 0.016}$ $\mathbf{0.542\pm 0.017}$ $\mathbf{0.604\pm 0.013}$ $\mathbf{0.658\pm 0.019}$ $\mathbf{0.544\pm 0.026}$ $\mathbf{0.660\pm 0.019}$
EEG + iEEG $0.514\pm 0.110$ $0.453\pm 0.123$ $0.521\pm 0.116$ $0.651\pm 0.013$ $0.534\pm 0.017$ $0.653\pm 0.013$
16 EEG-only $0.541\pm 0.048$ $0.485\pm 0.054$ $0.551\pm 0.051$ $0.662\pm 0.012$ $0.549\pm 0.017$ $0.664\pm 0.012$
EEG + iEEG $\mathbf{0.601\pm 0.008}$ $\mathbf{0.550\pm 0.009}$ $\mathbf{0.607\pm 0.009}$ $\mathbf{0.676\pm 0.003}$ $\mathbf{0.567\pm 0.004}$ $\mathbf{0.678\pm 0.004}$
@@
[…]
@@ D.5 Interpretation Results
Figure 14 shows a visualization of representation analysis on downstream datasets using UMAP (McInnes et al., 2018). We examine the embeddings obtained from the pretrained $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ model on the test set of speech, onset, and volume tasks in neuroprobe dataset without any finetuning. The results indicate that DIVER learned meaningful iEEG representations from pretraining, thereby capturing label-relevant structure in downstream datasets even in the absence of finetuning.
@@
To obtain an interpretable salience measure of the model, we applied attention rollout (Abnar and Zuidema, 2020), which aggregates attention weights across all attention layers. Although DIVER’s any-variate attention produces a high-dimensional matrix, the attention rollout yielded a rank-1 matrix across all subjects and sessions. We therefore used a single representative row, reflecting a uniform attention distribution across keys independent of query. Figure 15 shows task-dependent temporal and cortical attribution patterns derived from the attention rollout of $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$. Attention rollout weights were computed on test sets for four high AUROC scored Neuroprobe tasks (volume, pitch, GPT-2 surprisal, and word head position). We follow PopT’s method of mapping attention weights onto the cortical surface, normalizing within each subject and scaling by the test AUROC (Chau et al., 2025). The resulting weights were averaged across subjects, trials, and folds.
Across both auditory (volume, pitch) and language (GPT-2 surprisal, word head position) tasks, attention rollout exhibits two temporal peaks: an early window (0–300 ms) and a late window (800–1000 ms). The early peak coincides with word onset of each sample and engages core speech–language regions, including inferior temporal gyrus and Broca’s and Wernicke’s areas (Mesgarani et al., 2014). The late peak shows task-specific spatial patterns. For auditory tasks, attribution remains concentrated in primary auditory cortex (Heschl’s gyrus and superior temporal gyrus). This is consistent with the possible continued auditory processing of subsequent words within the 1s window. In contrast, language tasks show increased engagement of temporal-frontal regions in the later window. This aligns with the known post-auditory latency of higher language networks (Hickok and Poeppel, 2007; Rauschecker and Scott, 2009). Overall, these results indicate that DIVER attends to tokens in a temporally structured manner with task-relevant spatial attribution. Notably, the right middle occipital gyrus remains consistently salient for GPT-2 surprisal, suggesting a contribution of visual-context representations to surprisal estimation (Friederici, 2011; Dehaene and Cohen, 2011). Collectively, DIVER’s task-dependent attributions are consistent with established neural models of auditory and language processing (Hickok and Poeppel, 2007).
Figure 15: Scaled attention rollout weights mapped onto the cortical surface for four Neuroprobe downstream tasks. Each panel shows $\textsc{DIVER}_{\text{Tiny}/\text{I}/0.1s}$ scaled attention weights extracted via attention rollout and aggregated in 100 ms bins from 0 ms to 1 s relative to stimulus onset. For each task, the top row shows the left hemisphere and the bottom row shows the right hemisphere; brighter colors indicate larger scaled attention weights, which we interpret as greater model salience on the corresponding channel-temporal token for each task.
@@ E.1 Overview: Impact of Finetuning Methodology on EEG Downstream Task Performance
We conduct comprehensive experiments to investigate the impact of different finetuning methods on downstream task performance. Our analysis includes: (1) systematic exploration of MLP classifier depth for DIVER(appendix subsection E.2), (2) reproduction of CBraMod (Wang et al., 2024c) performance using publicly available code and weights(appendix subsection E.3), and (3) comparative analysis of various finetuning configurations across both models E.4. These experiments reveal that finetuning methodology significantly affects model performance, and optimal configurations vary across tasks and models. Importantly, when comparing DIVER and CBraMod under identical finetuning configurations (one-to-one comparison), DIVER demonstrates competitive or superior performance to CBraMod across multiple tasks, achieving overall state-of-the-art results despite CBraMod’s higher reported in-paper performance on some tasks.
@@ E.2 Impact of MLP Classifier Depth on DIVER Performance
Initially, we employed a linear classifier head for finetuning but observed suboptimal performance. Upon examining the publicly available code of CBraMod (Wang et al., 2024c), the previous state-of-the-art model, we found that it uses an MLP classifier head. To ensure fair comparison, we replaced the linear head with an MLP classifier, which resulted in substantial performance improvements for DIVER-1. This motivated us to conduct systematic experiments investigating how MLP depth affects performance. Table 23 compares the performance across the depths of MLP classifier for downstream EEG tasks. The original 3-layer MLP classifier is varied between 1 to 5 layers. For both FACED and PhysioNet-MI tasks, performance improved with the increase of MLP depth, achieving peak performance at 4-layer for FACED (balanced accuracy of 0.603) and 3-layer for PhysioNet-MI (balanced accuracy of 0.676). Beyond the optimal depth, we observed performance saturation or slight degradation, particularly notable in the 5-layer MLP for both datasets. These results indicate that a moderate depth (3-4 layers) suffices model’s effectiveness across different EEG downstream tasks.
@@ E.3 Challenges in Reproducing CBraMod Baseline Performance
The experimental results revealed substantial performance gaps on several tasks (Table 24). On MentalArithmetic(Zyma et al., 2019), reproduced accuracy (0.619) fell short of reported performance (0.726); on TUEV(Obeid and Picone, 2016), accuracy decreased from 0.667 to 0.605; on Mumtaz2016(Zyma et al., 2019), from 0.956 to 0.882. Only FACED(Chen et al., 2023) and PhysioNet-MI(Goldberger et al., 2000; Schalk et al., 2004) showed relatively successful reproduction. Notably, similar reproduction difficulties with CBraMod have been reported in recent work (Wang et al., 2025), which also observed performance gaps between reported and reproduced results on certain tasks.
@@
Table 24: Comparison of DIVER-1 EEG model with CBraMod. CBraMod (in paper) refers to the performance reported in the original CBraMod paper, while CBraMod (reproduction) represents our reproduction using the default configuration from their publicly released code and weights. We conduct systematic finetuning method experiments to ensure fair comparison, with detailed results provided in Table 25.
[…]
black
@@ E.4 Task-Specific Optimal Finetuning Configurations for CBraMod and DIVER
Table 25: One-to-One Comparison between CBraMod and DIVER using fixed finetuning method. Both models evaluated with identical finetuning configurations. CBraMod uses publicly released weights and code. DIVER model uses 12 layers with $d_{\text{model}}=512$ and were pretrained for 16 epochs. Results are reported as mean $\pm$ standard deviation across 5 random seeds.
@@
E.5 Linear Probing Analysis: Evidence of High-Quality Learned Representations
Linear probing provides a controlled method to evaluate the intrinsic quality of learned representations by eliminating confounding effects of different finetuning configurations (Radford et al., 2021; Chen et al., 2020). Recent electrophysiology foundation model research has similarly employed linear probing to assess representation quality (Anonymous, 2026). Given that optimal finetuning configurations vary across models and tasks (Section E.4), we conduct additional analysis using linear probing to assess the intrinsic quality of DIVER’s learned representations. Table 26 presents linear probing results across five downstream tasks, comparing DIVER against CBraMod (Wang et al., 2024c) and EEGPT (Wang et al., 2024b).
DIVER demonstrates strong linear probing performance across all tasks. On FACED, MentalArithmetic, and Mumtaz2016, DIVER achieves the best performance across all metrics. On TUEV, DIVER achieves the highest accuracy (0.559), while EEGPT shows stronger performance on kappa and F1 scores. On PhysioNet-MI, CBraMod achieves the best performance, though DIVER remains competitive as the second-best model.
These linear probing results indicate that DIVER learns high-quality embeddings that effectively capture task-relevant information even without tuning the backbone. This suggests that DIVER’s pretraining approach successfully learns generalizable representations of electrophysiological signals across diverse downstream applications.
Table 26: Linear Probing Performance Comparison. DIVER demonstrates superior linear probing performance on 4 out of 5 tasks, indicating high-quality learned representations. Bold indicates best performance; underline indicates second-best. Results are reported as mean $\pm$ standard deviation across 5 random seeds.
[…]
black
Models FACED (9-class)
ACC kappa F1
CBraMod $\underline{0.257\pm 0.010}$ $\underline{0.160\pm 0.011}$ $\underline{0.216\pm 0.012}$
EEGPT $186\pm 0.011 0.0$$084\pm 0.012$ $0.183\pm 0.016$
DIVER (Ours) $\mathbf{0.355\pm 0.005}$ $\mathbf{0.272\pm 0.005}$ $\mathbf{0.352\pm 0.004}$
Models PhysioNet-MI (4-class)
ACC kappa F1
CBraMod $\mathbf{0.545\pm 0.011}$ $\mathbf{0.393\pm 0.015}$ $\mathbf{0.546\pm 0.014}$
EEGPT $0.486\pm 0.010$ $0.290\pm 0.014$ $0.468\pm 0.009$
DIVER (Ours) $\underline{0.530\pm 0.011}$ $\underline{0.373\pm 0.015}$ $\underline{0.531\pm 0.011}$
Models MentalArithmetic (2-class)
ACC AUC-PR AUROC
CBraMod $0.515\pm 0.008$ $\underline{0.522\pm 0.017}$ $\underline{0.668\pm 0.008}$
EEGPT $\underline{0.535\pm 0.028}$ $0.335\pm 0.042$ $0.590\pm 0.045$
DIVER (Ours) $\mathbf{0.608\pm 0.037}$ $\mathbf{0.667\pm 0.021}$ $\mathbf{0.791\pm 0.011}$
Models Mumtaz2016 (2-class)
ACC AUC-PR AUROC
CBraMod $0.515\pm 0.003$ $0.962\pm 0.003$ $\underline{0.972\pm 0.002}$
EEGPT $0.897\pm 0.014$ $\underline{0.970\pm 0.006}$ $0.968\pm 0.006$
DIVER (Ours) $\mathbf{0.944\pm 0.003}$ $\mathbf{0.990\pm 0.000}$ $\mathbf{0.990\pm 0.000}$
Models TUEV (6-class)
ACC kappa F1
CBraMod $0.314\pm 0.011$ $0.307\pm 0.014$ $0.574\pm 0.012$
EEGPT $\underline{0.446\pm 0.055}$ $\mathbf{0.431\pm 0.059}$ $\mathbf{0.710\pm 0.029}$
DIVER (Ours) $\mathbf{0.559\pm 0.025}$ $\underline{0.397\pm 0.037}$ $\underline{0.624\pm 0.041}$
@@ F.1 Pretraining Dataset Description
  • AJILE12 (Anootated Joints in Long-term Electrocorticography) (Peterson et al., 2022): An ECoG dataset from 12 epilepsy patients, recorded semi-continuously over 55 days. Signals were collected from $\geq$ 64 electrodes at 1 kHz sampling rate and paired with synchronized video-vased 3D human pose estimation and annotated wrist-movement events.
@@
  • TUEG (Temple University Hospital EEG Corpus) (Obeid and Picone, 2016): A large-scale clinical EEG dataset comprising 16,986 recording sessions from 10,874 subjects with heterogeneous diagnoses. EEG signals were recorded using 20–31+ channels, predominanty at sampling rate between 250–512 Hz, and are linked with de-identified clinical reports.
  • HBN-EEG (Healthy Brain Network) (Shirazi et al., 2024): An developmental EEG dataset from 2,782 participants aged 5–21. Each participant underwent approximately 60 minutes of high-density (128-channel) EEG and eye-tracking recordings across six distinct tasks, including resting-state and movie watching.
  • NCHSDB (Nationwide Children’s Hospital Sleep DataBank) (Lee et al., 2022): An pediatric sleep EEG dataset of 3,673 patients. Each record includes 8–12 hours of EEG data (26–29 channels, sampled at 256–512 Hz) manually scored for sleep stages and events. While the dataset contains multimodal PSG signals (EOG, EMG, ECG, respiration, etc.), we used only the EEG channels.
  • PEERS (Penn Electrophysiology of Encoding and Retrieval Study) (Kahana et al., 2023): An EEG dataset from 364 subjects who participated in multiple sessions of free recall, recognitionm and distractor tasks. EEG signals were recorded with 125 channels at 500 Hz sampling rate.